# A Curriculum Learning Approach to Reinforcement Learning: Leveraging RAG for Multimodal Question Answering

### v1

@@ 2.1. Visual Question Answering
The advent of VLLMs has revolutionized the VQA field. Early notable works include ViLBERT (Lu et al., 2019) and VL-BERT (Su et al., 2019), which extended BERT architecture to handle multi-modal inputs. LXMERT (Tan and Bansal, 2019) proposed cross-modality pre-training to learn joint visual and language representations. More recent approaches shifted toward unified Transformer frameworks and contrastive learning methods exemplified by CLIP (Radford et al., 2021), which demonstrated powerful zero-shot learning capabilities. Contemporary VLLMs employ various pre-training strategies including masked modeling techniques and generative approaches. The latest trend involves fine-tuning existing large language models (LLMs) like Llama (Meta, 2023), Qwen (Bai et al., 2023), and Falcon using adapters that map visual tokens to language tokens, significantly reducing training resources while maintaining state-of-the-art performance across various benchmarks (Liu et al., 2023; Zhu et al., 2023).
@@ 3.1. Supervised Fine-tuning
During supervised fine-tuning, we focus on enhancing the model’s ability to generate accurate chain-of-thought (CoT) style answers in the required <think></think> and <answer></answer> format, rather than directly controlling hallucinations at this stage. To construct high-quality SFT training data, we distill both answers and detailed reasoning processes from stronger models such as GPT-4.1 (Achiam et al., 2023). In our prompting strategy, GPT-4.1 is provided with the ground-truth answer and is instructed to decide whether the question is answerable. If so, it generates both the answer and a corresponding CoT; if not, it returns an explicit refusal and the reason for being unable to answer. This ensures that the SFT data reflects both high-quality reasoning and the ability to abstain when appropriate.
@@
In addition, for the integrated auxiliary search tools such as image search and web search, the model needs to acquire the ability to invoke these tools. Specifically, we leverage more powerful models (such as GPT-4.1) to generate explicit CoT outputs for tool usage, and use these data to fine-tune the Llama Vision model. This enables the model to, given an image and a question, recognize what knowledge it lacks and determine which search tool to use and what search query to issue. Concretely, if the model needs to use image search, it should output the coordinates of the relevant sub-region of the original image; if web search is required, it should output one or more text queries. Our experiments show that the use of image search is not effective, and this will be discussed in detail in subsequent sections. For the additional knowledge retrieved via search, we concatenate it with the original question to help the model generate the final answer.
@@ 3.3. Retrieval-Augmented Generation
To better leverage web search, we developed a dedicated retriever. Specifically, for a given query, the retriever uses web search to obtain k web pages, extracts the main content using BeautifulSoup, and segments both the main text and summary. We perform multi-stage retrieval: initial coarse ranking is done using a combination of bge-large-en-v1.5 embeddings, BM25, and TF-IDF, followed by re-ranking with bge-reranker-v2-m3 to select the top-k results.
@@
  • Multi-turn augmentation (Task 3): The knowledge augmentation process in Task 3 is similar to that of Task 2, with the main difference being that Task 3 involves multi-turn dialogue, as opposed to the single-turn setting in Task 2. In each turn, the model can decide whether additional external knowledge is needed for the current question.
@@ 4.1.1. Dataset
The CRAG-MM dataset (Yang et al., 2024) is designed to benchmark multi-modal, retrieval-augmented question answering systems and consists of three main components: an image set, a collection of question-answer (QA) pairs, and retrieval contents.
@@ 4.1.2. Metrics
For multi-turn QA, following (Xu et al., 2023), a conversation is terminated if two consecutive answers are labeled as ”Incorrect”, and all subsequent answers are considered ”Missing”. The average score across all multi-turn conversations is then reported as the final truthfulness score.

### v2

@@ 2.1. Visual Question Answering
The advent of VLLMs has revolutionized the VQA field. Early notable works include ViLBERT (Lu et al., 2019) and VL-BERT (Su et al., 2019), which extended BERT architecture to handle multi-modal inputs. LXMERT (Tan and Bansal, 2019) proposed cross-modality pre-training to learn joint visual and language representations. More recent approaches shifted toward unified Transformer frameworks and contrastive learning methods exemplified by CLIP (Radford et al., 2021), which demonstrated powerful zero-shot learning capabilities. Contemporary VLLMs employ various pre-training strategies including masked modeling techniques and generative approaches. The latest trend involves fine-tuning existing large language models (LLMs) like Llama (Meta, 2023), Qwen (Bai and others, 2023), and Falcon using adapters that map visual tokens to language tokens, significantly reducing training resources while maintaining state-of-the-art performance across various benchmarks (Liu and others, 2023; Zhu et al., 2023).
@@ 3.1. Supervised Fine-tuning
During supervised fine-tuning, we focus on enhancing the model’s ability to generate accurate chain-of-thought (CoT) style answers in the required <think></think> and <answer></answer> format, rather than directly controlling hallucinations at this stage. To construct high-quality SFT training data, we distill both answers and detailed reasoning processes from stronger models such as GPT-4.1 (Achiam et al., 2023). In our prompting strategy, GPT-4.1 is provided with the ground-truth answer and is instructed to decide whether the question is answerable. If so, it generates both the answer and a corresponding CoT; if not, it returns an explicit refusal and the reason for being unable to answer. This ensures that the SFT data reflects both high-quality reasoning and the ability to abstain when appropriate. We use the following prompt to generate responses with GPT-4.1:
[{"role": "user", "content": "You are an intelligent assistant
that can answer users’ questions based on the pictures they provide.
You need to put your thinking/reasoning  process into <think>xxx
</think> and put the answer or conclusion into <answer>xxx</answer>.
Please avoid wrong answers and hallucinations. When you can get
a high-confidence answer, <answer>xxx </answer> is your answer;
When your current knowledge is not enough for you to answer the
question, <answer>xxx</answer> should be  <answer>I don’t know,
please add xxx information</answer>; don’t give ambigious answer.
Your response should strictly follow the format of <think>xxx
</think>, <answer>xxx</answer>. "}
{"role": "user", "content":"Please answer the following
question based on the provided image: Please answer the following
question based on the provided image: {query_str}"},
{"role": "user", "content": "The answer is: {answer_str}"},
{"role": "user", "content": [{"type": "image"}]}]
It is important to note that during the SFT of the Llama model, the answer is not provided.
@@
In addition, for the integrated auxiliary search tools such as image search and web search, the model needs to acquire the ability to invoke these tools. Specifically, we leverage more powerful models (such as GPT-4.1) to generate explicit CoT outputs for tool usage, and use these data to fine-tune the Llama Vision model. This enables the model to, given an image and a question, recognize what knowledge it lacks and determine which search tool to use and what search query to issue. Concretely, if the model needs to use image search, it should output the coordinates of the relevant sub-region of the original image; if web search is required, it should output one or more text queries. Our experiments show that the use of image search is not effective, and this will be discussed in detail in subsequent sections. For the additional knowledge retrieved via search, we concatenate it with the original question to help the model generate the final answer. We use the following prompt to generate querys:
[{"role": "user", "content": "You are a RAG search assistant.
For the given question and picture, understand and analyze the
question, abstract the specific question, and analyze the purpose
of the question.Please summarize a query from the question
( the query will be used to search for relevant information
in the text rag later) to help the subsequent small model answer
the current question. Your query should be detailed and concise,
and cover the key information to solve the current problem, such
as: <BMW i5 2025 price><Audi q5l engine><Dragon fruit China distri-
bution><Azalea flowering period><White House construction date>
<Eiffel Tower designer> etc. Please write the reasoning process
in <think></think> and the query in <search></search>.
-- Your reasoning process should be detailed and step-by-step.
-- When the current problem is simple recognition and simple calcu-
lation, no additional search information is required, and your
query can be empty. That is, <think>reasoning</think><search>
</search>, it must be strictly  in this form. In most cases, you
should use query first, because this information can help the
downstream small model answer questions and effectively avoid
the illusion of the small model."},
{"role": "user", "content":"The question is: {query_str}"},
{"role": "user", "content": [{"type": "image"}]}]
This is the prompt to generate responses using retrieval informaton:
[{"role": "user", "content": "You are an intelligent assistant
and can answer users’ questions based on the pictures provided
by users. Please avoid giving wrong answers and having halluci-
nations.When you can obtain a highly reliable answer, return the
answer directly; When your existing knowledge is insufficient to
answer the question, your answer should be ’I don’t know.’’ For
Yes/No questions, the answers should start with ’Yes’ or ’NO’, and
no ambiguous answers should be given. In addition, the web search
content will also be provided for your reference. Please carefully
verify whether the provided search information is reliable enough
to help you answer correctly. The search information is
{search_result}."}
{"role": "user", "content":"The question is: {query_str}"},
{"role": "user", "content": [{"type": "image"}]}]
@@ 3.2. Reinforcement Learning with Curriculum Learning
During the GRPO phase, while focusing on enhancing answer accuracy, there is an observed trade-off where the model’s ability to generate queries might slightly diminish. This is acceptable because our primary goal is to improve the accuracy of answers and effectively control hallucinations. Despite this trade-off, the model retains enough query generation capability to support complex question answering.
@@ 3.3. Retrieval-Augmented Generation
To better leverage web search, we developed a dedicated retriever. Specifically, for a given query, the retriever uses web search to obtain k web pages, extracts the main content using BeautifulSoup, and segments both the main text and summary. We perform multi-stage retrieval: initial coarse ranking is done using a combination of bge-large-en-v1.5 embeddings, BM25, and TF-IDF, followed by re-ranking with bge-reranker-v2-m3 to select the top-k results. The actual k value used for both web retrieval and the final answer selection is 5.
@@
  • Multi-turn augmentation (Task 3): The knowledge augmentation process in Task 3 is similar to that of Task 2, with the main difference being that Task 3 involves multi-turn dialogue, as opposed to the single-turn setting in Task 2. Thus, we provide historical questions and answers as part of the reference information. The VLLM itself determines whether and how to use this historical data, allowing it to leverage past interactions to potentially enhance its understanding and response quality.
@@ 4.1.1. Dataset
The CRAG-MM dataset (Wang et al., 2025) is designed to benchmark multi-modal, retrieval-augmented question answering systems and consists of three main components: an image set, a collection of question-answer (QA) pairs, and retrieval contents.
@@ 4.1.2. Metrics
For multi-turn QA, following (Xu and others, 2023), a conversation is terminated if two consecutive answers are labeled as ”Incorrect”, and all subsequent answers are considered ”Missing”. The average score across all multi-turn conversations is then reported as the final truthfulness score.